A CaCO3-MgCO3 solid solution contains 3 weight percentage Mg. What is the mole fraction of MgCO3 in the solid?


1
Expert's answer
2021-08-09T07:46:57-0400

Solution:

The molar mass of CaCO3-MgCO3 is 184 g/mol.

Suppose that 184 g (or 1 mol) of a CaCO3-MgCO3 solid solution are given.

Therefore,

m(Mg) = [w(Mg) × m(CaCO3-MgCO3 solid solution)] / 100% = (3% × 184 g) / 100% = 5.52 g Mg


The molar mass of Mg is 24 g/mol.

Therefore,

(5.52 g Mg) × (1 mol Mg / 24 g Mg) = 0.23 mol Mg


MgCO3 → Mg

According to the equation above: n(Mg) = n(MgCO3) = 0.23 mol


Thus,

χ(MgCO3) = n(MgCO3) / n(CaCO3-MgCO3)

χ(MgCO3) = (0.23 mol) / (1 mol) = 0.23

χ(MgCO3) = 0.23


Answer: The mole fraction of MgCO3 in the solid is 0.23